OpenStudy (anonymous):

A plane flying horizontally at an altitude of 1 mi and a speed of 530 mi/h passes directly over a radar station. Find the rate at which the distance from the plane to the station is increasing when it is 5 mi away from the station. (Round your answer to the nearest whole number.)

OpenStudy (aum):



OpenStudy (anonymous):

thats confusing

OpenStudy (aum):

\[ s^2 = x^2 + 1 \\ \text{Differentiate with respect to t:} \\ 2s\frac{ds}{dt} = 2x\frac{dx}{dt} \\ 2s\frac{ds}{dt} = 2\sqrt{s^2-1}\frac{dx}{dt} \\ s = 5~\text{miles}; \frac{dx}{dt} = 530 ~\text{mph} \\ 2 * 5 * \frac{ds}{dt} = 2 * \sqrt{25-1} * 530 \\ \text{Solve for }\frac{ds}{dt} \]

OpenStudy (anonymous):

i got the wrong answer

OpenStudy (aum):

What answer did you get?

OpenStudy (anonymous):

nevermind i got it